OpenStudy (anonymous):

Evaluate the improper intergral:

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x ^{\frac{ -1 }{ 2 }}dx\]

OpenStudy (reemii):

the method is to compute first \(\int_\varepsilon^1 x^{-1/2} \,dx\). This will be an expression involving the \(\varepsilon\). then compute the limit for of this function of \(\varepsilon\) when \(\varepsilon\to0+\).

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} xe ^{-x ^{2}}dx \] @reemii !!! Evaluate the improper integral

OpenStudy (reemii):

isn't the definition: this integral exists if both improper integrals exist ? \(\int_{-\infty}^0 f(x)dx\) and \(\int_0^{\infty} f(x)dx\)

OpenStudy (reemii):

yes it's that. So, concentrating on \(\int_0^\infty x e^{-x^2}dx\), .. where do you think the \(x\) is coming from?

OpenStudy (anonymous):

I have no idea what ur asking

OpenStudy (reemii):

oh, I mean, if you compute the derivative of \(e^{-x^2}\) , you'll obtain \(e^{-x^2}\times x\times C\). So, .. integrating this function is very easy.

OpenStudy (reemii):

C=-2

OpenStudy (anonymous):

I have an answere sheet and it says the answere is 0....

OpenStudy (reemii):

yes it's correct. the integral on the right is \([-\frac12 e^{-x^2}]_0^\infty = 0 - (-1)\). On the "left", it's equal to -1. therefore the complete integral is -1+1= 0.

OpenStudy (anonymous):

Ohhh. ok!

OpenStudy (reemii):

oops, rather: -1/2 + 1/2 = 0 ^^'